OpenStudy (anonymous):

Addison has been told that her mean average on six homework assignments in her history class is 85. She can find only five of the six assignments, which have scores of 91,72,96,88, and 75. Find the score on the lost homework. Please help me show how to solve this using the mean "x" for the missing score

OpenStudy (kirbykirby):

The mean is \[\frac{x_1+x_2+x_3+x_4+x_5+x_6}{6}=\bar{x} \] You have the following: \[\frac{ 91+72+96+88+ 75+x_6}{6}=85 \] Can you solve for \(x_6\)?

OpenStudy (anonymous):

88

OpenStudy (kirbykirby):

yes :)